OpenStudy (anonymous):

What if it only says "The average of 3 test scores if the first two scores are 78 and 82 s=3rd test score "???

OpenStudy (anonymous):

(78+82)/2=80 (3rd test score) now avrg of 78,82,80 gives the answer

OpenStudy (anonymous):

thanks :D

OpenStudy (whpalmer4):

@Newtons_apple you can't really do that. averaging has to be done over all the scores at once, not averaging some scores with the average of the others. consider averaging 4, 6, 8 and 10. 4+6+8+10 = 28 28/4 = 7, so 7 is the average of those 4. If you however average the first 3 (4+6+8=18, 18/3 = 6) and then average 6 and 10 together, you get 8, not 7.

OpenStudy (anonymous):

,m